OpenStudy (anonymous):

**Can somebody please check my work** I have to build a landscape wall that’s 2ftx20ft using either vineyard stones, cottage stones, or manor stones. It takes 4.25 vineyard stones, 3 cottage stones, and 1.5 manor stone to equal 1 square foot. The vineyard stones are $0.89 each, cottage stones are $1.18 each, and manor stones are $2.67 each. 1. How many stones would you need to build the wall for each option? 2. How much would each option cost? 3. Which option is the best deal? *2*20=40ft^2=area of landscape wall* #1 (Vineyard stones) 4.25*40=170 #1 (Cottage stones) 3*40=120

OpenStudy (anonymous):

#1 (Manor stones) 1.5*40=60 #2 (Vineyard stones) 0.89*170=151.30 #2 (Cottage stones) 1.18*120=141.60 #2 (Manor stones) 2.67*60=160.20 #3 Cottage stones are the best deal.